Required information

[The following information applies to the questions displayed below.]

Bargain Rental Car offers rental cars in an off-airport location near a major tourist destination in California. Management would like to better understand the variable and fixed portions of it car washing costs. The company operates its own car wash facility in which each rental car that is returned is thoroughly cleaned before being released for rental to another customer. Management believes that the variable portion of its car washing costs relates to the number of rental returns. Accordingly, the following data have been compiled:

Month Rental Returns Car Wash Costs
January 2,400 $ 11,400
February 2,500 $ 13,300
March 2,800 $ 12,200
April 3,100 $ 14,900
May 3,700 $ 16,600
June 5,100 $ 24,100
July 5,600 $ 22,600
August 5,600 $ 23,200
September 4,800 $ 23,200
October 4,100 $ 22,300
November 2,200 $ 11,100
December 3,000 $ 15,600

2. Using least-squares regression, estimate the variable cost per rental return and the monthly fixed cost incurred to wash cars. (Round Fixed cost to the nearest whole dollar amount and the Variable cost per unit to 2 decimal places.)

fixed cost per month:

variable cost per rental return:
Answer #1
Month Rental Returns(X) Total Cost(Y)(in $) Square of X X*Y(in $)
Jan                                    2,400                   11,400         57,60,000               2,73,60,000
Feb                                    2,500                   13,300         62,50,000               3,32,50,000
Mar                                    2,800                   12,200         78,40,000               3,41,60,000
Apr                                    3,100                   14,900         96,10,000               4,61,90,000
May                                    3,700                   16,600      1,36,90,000               6,14,20,000
Jun                                    5,100                   24,100      2,60,10,000             12,29,10,000
Jul                                    5,600                   22,600      3,13,60,000             12,65,60,000
Aug                                    5,600                   23,200      3,13,60,000             12,99,20,000
Sep                                    4,800                   23,200      2,30,40,000             11,13,60,000
Oct                                    4,100                   22,300      1,68,10,000               9,14,30,000
Nov                                    2,200                   11,100         48,40,000               2,44,20,000
Dec                                    3,000                   15,600         90,00,000               4,68,00,000
Total                                  44,900                2,10,500    18,55,70,000             85,57,80,000
Unit Variable Cost (b) = nΣxy − (Σx)(Σy) /nΣx^2 − (Σx)^2
Total Fixed Cost (a) = Σy − bΣx / n
So here n =12 ,Σx =44.900, Σy =210,500, Σx^2=18,55,70,000 ,Σxy =85,57,80,000
Variable cost per unit(b) =12*85,57,80,000 - (44,900*210,500) / 12*18,55,70,000 - 44,900^2
Variable cost per unit(b) =10,26,93,60,000 - 9,45,14,40,000 / 2,03,18,40,000 - 2,01,60,10,000
Variable cost per unit(b) =81,79,10,000 / 21,08,30,000 =$3.88
Total Fixed cost =210,500 - ($3.88*44,900) / 12
Total Fixed cost =(210,500 - 174,212) / 12
Total Fixed cost =$36,288 /12 =$3,024
The cost volume formula is:
y =$3,024 + $3.88X